OpenStudy (anonymous):

Which of the following is an identity? A. sin x - cos x + 1 = tan x B. (1 - 2sin^2 x)csc^2 x = 4cos^2 x - 2 C. sin^2 x cot^2 x + cos^2 x tan^2 x = 1 D. tan^2 x + cot^2 x = 1 ***i think it's either A or C, but I can't figure out which one it is.... what do u think? @terenzreignz ? :)

terenzreignz (terenzreignz):

Tricky... hold on a sec...

OpenStudy (anonymous):

kk:)

terenzreignz (terenzreignz):

okay.... which one is your guess? come now...

OpenStudy (anonymous):

so i think it's either A or C, but as i'm looking at it now, i'm leaning a lot more towards A.. is that a good feeling that I'm feeling when i'm feeling that it should be A? ;p hehe or is it just BAD? hahahaha

terenzreignz (terenzreignz):

Well, too bad... \[\huge \sin^2(x)\cot^2(x)=\sin^2(x)\cdot\frac{\cos^2(x)}{\sin^2(x) }=\color{red}? \]

OpenStudy (anonymous):

aww gosh darn :( it was a bad feeling then.... :( darn. it equals cos^2 x? cuz the sin^2 x cancels out right?

terenzreignz (terenzreignz):

Yes, precisely :) Also, similarly... \[\huge \cos^2(x)\tan^2(x)=\cos^2(x)\cdot\frac{\sin^2(x)}{\cos^2(x) }=\color{green}?\]

OpenStudy (anonymous):

sin^2 x?

terenzreignz (terenzreignz):

Yes... so in fact, C becomes... \[\large \color{red}{\sin^2(x)\cot^2(x)}+\color{green}{\cos^2(x)\tan^2(x)} = 1\] May be simplified into...?

OpenStudy (anonymous):

does it simplify to this? sin^2 x cot^2 x + cos^2 x tan^2 x = 1

terenzreignz (terenzreignz):

Sorry, I meant the left side simplifies into...?

terenzreignz (terenzreignz):

Given those two preliminary questions I asked you (LOL)

OpenStudy (anonymous):

wait what? hahaha confused :/

OpenStudy (anonymous):

what am i simplifying? :/

terenzreignz (terenzreignz):

\[\huge \color{red}{\sin^2(x)\cot^2(x)}=\color{blue}{\cos^2(x)}\\ \huge \color{green}{\cos^2(x)\tan^2(x)}=\color{blue}{\sin^2(x)}\] Now please simplify the left side...\[\large \color{red}{\sin^2(x)\cot^2(x)}+\color{green}{\cos^2(x)\tan^2(x)} = 1\]

OpenStudy (anonymous):

ohhh okay... so cos^2 x + sin^2 x ?

terenzreignz (terenzreignz):

Yup... so it becomes \[\huge \color{blue}{\cos^2(x)+\sin^2(x)}=1\] Which is always true, right?

OpenStudy (anonymous):

yes, at least i believe so! sooo what does that mean? :O is the answer C then? :)

terenzreignz (terenzreignz):

Yes. ;)

terenzreignz (terenzreignz):

Did your score go up a point? :/

OpenStudy (anonymous):

yay!! :) haha stupid gut... and my stupid feelings :P i had a feeling it was either A or C.. then i had a feeling it was A, and boy was that feeling wrong hahaha :P oh well... i'll give my feelings a second chance ;) thanks!!! :D my score? :/

OpenStudy (anonymous):

idk what you mean haha? what score? :/

terenzreignz (terenzreignz):

Never mind :) Quiero ver tu ultima pregunta :)

OpenStudy (anonymous):

haha coming right up! :)

OpenStudy (anonymous):

haha coming right up! :)